### v1

@@ Abstract
Climate change exacerbates extreme weather events like heavy rainfall and flooding. As these events cause severe losses of property and lives, accurate high-resolution simulation of precipitation is imperative. However, existing Earth System Models (ESMs) struggle with resolving small-scale dynamics and suffer from biases, especially for extreme events. Traditional statistical bias correction and downscaling methods fall short in improving spatial structure, while recent deep learning methods lack controllability over the output and suffer from unstable training. Here, we propose a novel machine learning framework for simultaneous bias correction and downscaling. We train a generative diffusion model in a supervised way purely on observational data. We map observational and ESM data to a shared embedding space, where both are unbiased towards each other and train a conditional diffusion model to reverse the mapping. Our method can be used to correct any ESM field, as the training is independent of the ESM. Our approach ensures statistical fidelity, preserves large-scale spatial patterns and outperforms existing methods especially regarding extreme events and small-scale spatial features that are crucial for impact assessments.
keywords:
precipitation simulation, statistical downscaling, bias correction, conditional diffusion models, extreme weather events, Earth system modelling
@@ 1 Introduction
With global warming, we anticipate more intense rainfall events and associated natural hazards, e.g. in terms of floods and landslides, in many regions of the world [16]. Understanding and accurately simulating precipitation is particularly important for adaptation planning and, hence, for mitigating damages and reducing risks associated with climate change. Earth System Models (ESMs) play a crucial role in simulating precipitation patterns for both historical and future scenarios. However, these simulations are computationally extremely demanding, primarily because they require solving complex partial differential equations. To manage the computational load, ESMs resort to approximate solutions on discretized grids with coarse spatial resolution (typically around 100 km). The consequence is that these models do not resolve small-scale dynamics such as many of the processes relevant to precipitation generation. This leads to considerable biases in ESM fields compared to observations. Moreover, the coarse spatial resolution prevents accurate projections of localized precipitation extremes. The precipitation fields simulated by ESMs can, therefore, not be directly used for impact assessments [28] and especially tasks like water resource and flood management, which demand precise spatial data at high resolution [7].
@@
Statistical bias correction methods can be used as a post-processing to adjust statistical biases. Quantile mapping (QM) is the most common method for improving the statistics of ESM precipitation fields [24, 6, 1]. QM reduces the bias using a mapping that, locally at each grid cell, aligns the estimated cumulative distribution of the model output with the observed precipitation patterns over a reference time period. While QM is effective in correcting distributions of single grid cells, it falls short in improving the spatial structure and patterns of precipitation simulations [10]. Visual inspection shows that ESM precipitation remains too smooth compared to observational data after applying quantile mapping (Fig. 3).
@@
To address these problems, deep learning methods have recently been introduced [11, 18, 4, 10]. In these approaches, the statistical relationships between model simulations and observational data are learned implicitly. A general constraint when using machine learning methods for bias correction is that individual samples of observational and Earth System Model data are always unpaired. A sample, in this context, is a specific weather situation at a specific point in time. The reason for this lack of pairs is that simulations, even with very similar initial conditions, diverge already after a short period of time due to the chaotic nature of the underlying atmospheric dynamics. Currently, one can, therefore, not utilize the wide range of supervised ML techniques that have shown great success in various disciplines in recent years and the available options are consequently restricted to self- and unsupervised machine learning methods. Recent studies [11, 18, 4, 10] applied generative adversarial networks (GANs [5]) and specifically cycleGANs [29] to improve upon existing bias correction techniques. A major limitation of GAN-based approaches is that stability and convergence of the training process are hard to control and that it is challenging to find metrics that indicate training convergence. Also, GANs often suffer from mode collapse, where only a part of the target probability distribution is approximated by the GAN.
@@
As noted above, the low spatial resolution of ESM fields prevents local risk and impact assessment, which necessitates the additional use of downscaling methods. In line with the climate literature, we refer to increasing the spatial resolution as downscaling throughout our manuscript, although we are aware that, especially in the machine learning literature, the term upsampling is more prevalent. We use the term downscaling only when we want to increase the information content in an image as well as the number of pixels. When we refer to upsampling (downsampling), we only mean an increase (decrease) in the number of pixels. Statistical downscaling aims to learn a transformation from the low resolution ESM fields to high resolution observations. Recent developments lean towards using machine learning methods for this task [19, 14]. The potential for machine learning based downscaling methods was shown already in [25]. The work presented a downscaling approach built on purely convolutional networks. More recent work [26] introduced a framework for downscaling and bias correction, combining a diffusion model for downscaling and a model based on optimal transport for bias correction [2]. Optimal transport learns a map between two data distributions in an unsupervised setting. The downside of their framework is its computational expense. The effectiveness also has not been demonstrated on real observational and ESM fields yet.
@@
Diffusion models (DMs) have recently emerged as the state-of-the-art ML approach for conditional image generation [22, 20, 23] and image-to-image translation [21], mostly outperforming GANs across different tasks. Diffusion models (Fig. 1 and S1) avoid the common issues present with GANs in exchange for slower speed at inference. A diffusion model consists of a forward and a backward process. During the forward process, noise is added to an image in subsequent steps to gradually remove its content. The amount of noise added follows a predefined equation. During the backward process, a neural network is trained to reverse each of these individual noising steps to recover the original image. The trained diffusion model can generate an image of the training data distribution, given a noise image as input.
We present here a framework based on state-of-the-art diffusion models that allows to perform both bias correction and downscaling with one single network. This network can, in principle, be any supervised machine learning method. We use a single conditional diffusion model (Fig. 1 and Fig. S1) to correct low resolution (LR) ESM fields towards high resolution (HR) observational data (OBS). In general, our task of bias correction and downscaling can be seen as taking a field from a distribution $p(ESM)$ and transforming it into a field from a conditional distribution $p(OBS|ESM)$. We cannot directly learn such a downscaling and bias correction transformation because ESM and OBS are unpaired. Our framework, therefore, introduces transformations $f$ and $g$ that map ESM and OBS data to a shared embedding space (see Methods and Fig. 1A). On this shared embedding space, we can train a conditional diffusion model to approximate the inverse of $f$. We are then able to bias correct and downscale ESM fields by first mapping them to the shared embedding space and then projecting them to the OBS dataset with the trained diffusion model (Fig. 1B and Fig. 1C). This framework offers great flexibility as it can be applied to any ESM.
@@
Figure 1: (A) Schematic overview of our approach. Bias correction and downscaling can be formulated as a mapping $\omega$ from the $ESM$ data space to the data space of observations ($OBS$) used for training. We first map both datasets to a shared embedding space and then learn the inverse of the mapping $f$ with a diffusion model $DM$. We achieve a correction of the ESM data by applying $DM\circ g$. (B) Our framework allows to train a single model for bias correction and downscaling in a supervised way despite the unpaired nature of observations (OBS) and Earth system model (ESM) fields. We construct functions $f,g$ that map $OBS\in\mathbf{V^{obs}}$ and $ESM\in\mathbf{V^{esm}}$ fields to a shared embedding space $\mathbf{V^{emb}}$. By inverting $f$, we can rewrite $\omega$ as $\omega=f^{-1}\circ g$. We learn the inverse $f^{-1}$ with a conditional diffusion model $DM$. This model is trained (blue arrow) on pairs to approximate the map from $f(OBS)$ to $OBS$ (see Methods for details). Because $f(OBS)$ and $g(ESM)$ share the embedding space (and are identically distributed by construction), we can evaluate (green arrow) the conditional diffusion model $DM$ on the embedded ESM data $g(ESM)$ and thereby approximate the bias correction and downscaling function $\omega=f^{-1}\circ g\approx DM\circ g$.
@@
(C) Left: Training process of the conditional diffusion model $DM\approx f^{-1}$. Note that the individual samples of the input $OBS$ and their embeddings $f(OBS)$, as well as the embeddings $f(OBS)$ and the output of $DM\approx f^{-1}$ are paired, respectively. Right: Inference process of $DM\approx f^{-1}$. In this case, the individual samples of the input $ESM$, their embeddings $g(ESM)$, and the output of $DM\approx f^{-1}$ are paired, respectively. It is not necessary for the training embedding samples to be paired with the inference embedding samples. See Fig. S1 for details.
@@ 2.1 Embedding evaluation
The transformations $f,g$ are chosen such that they map the observational ($OBS$) and model ($ESM$) data to a common embedding space $\mathbf{V^{emb}}$, where all samples are identically distributed. For constructing $f$ and $g$ we need $f(ERA5)$ and $g(GFDL)$ to be unbiased with respect to each other. The transformations need to be chosen such that the embedded data share the same distribution and the same power spectral density (PSD). We assess if they are statistically unbiased towards each other by analyzing their histograms latitude / longitude profiles, as well as their spatial spectra. Figure S2 shows that $f(ERA5)$ and $g(GFDL)$ have the same spatial distribution (Fig. S2A) with minor differences in temporal statistics shown by the histogram (Fig. S2B) and latitudinal/ longitudinal profiles (Fig. S2C and Fig. S2D).
Figure 2: Power spectral densities (PSDs) for different choices noising scale of the diffusion model. The nosing scale $s$ (dashed line) is a hyperparameter that can be chosen depending on the ESM and observational datasets, as well as on the specific task. For the maximal choice of $s$ (A) all information in the observations (ERA5) and model simulations (GFDL) is noised and thereby destroyed. Conditioning on pure noise makes the task equivalent to unconditional image generation. The diffusion model will learn to generate observational fields with no relation to the ESM fields. When $s$ is chosen to be minimal, there will be no noising and the conditional generation will directly replicate the condition, i.e. the ESM field. In (B) we chose $s$ as the point where the PSDs of the observational and simulated datasets intersect. We then apply sufficiently many forward diffusion noising steps to both datasets, destroying small-scale structure until they agree in the PSD. We call scales smaller than $s$ small scales and scales larger than $s$ large scales. In (C) and (D), the effects of choosing a smaller noise scale $s$ are shown. Prior knowledge about the ESM or its accuracy can also guide the choice of $s$.
@@
The individual operations that make up the transformations $f$ and $g$ do not change the large-scale patterns of their respective inputs, as desired for a valid bias correction. The goal of the downscaling and bias correction $\omega$ (Fig. 1) is to rely on the unbiased large-scale patterns of the $ESM$ and correct statistics, as well as small-scale patterns. The transformation $g$ conserves the unbiased information from the ESM by construction. Therefore, we want the diffusion model, approximating $f^{-1}$, to also preserve unbiased information.
@@
The extreme case of erasing all detail with large amounts of noise leads to learning the unconditional distribution $p(ERA5)$, which is then not a correction of $GFDL$ but a generative emulation of the ERA5 reanalysis data. We test this by adding the same amount of noise to the output of our diffusion model that was added to create $g(GFDL)$. This ensures that both the downscaled and bias corrected fields, as well as the original GFDL fields, lack the small-scale details up to the same point.
@@
To verify that large-scale patterns are preserved by the diffusion model, we compute image similarity metrics between the low pass filtered version of the input of the diffusion model (embedded ERA5 data $f(ERA5)$) and the low pass filtered output of the diffusion model $DM(f(ERA5))$. The output of the low pass filter leaves the large-scale features unchanged. The comparison yields an average structural similarity index (SSIM [27]) value of 0.8 and a Pearson correlation coefficient of 0.9 for the validation dataset. This verifies that large-scale patterns are well preserved by the diffusion model.
@@
Our diffusion model is able to reconstruct high-resolution fields following the ERA5 distribution from embedded ERA5 fields $f(ERA5)$, with only minor discrepancies in small-scale patterns (Fig. S2A). A comparison between the mean absolute spatial-temporal difference between GFDL and ERA5 fields yields a bias of 0.3 mm/d. The downscaling and bias correction of our diffusion model reduces this bias to 0.1 mm/d.
Our diffusion model thus approximates $f^{-1}$ well and we successfully created a shared embedding space in which $f(ERA5)$ and $g(GFDL)$ are identically distributed.

### v2

@@ Abstract
Climate change exacerbates extreme weather events like heavy rainfall and flooding. As these events cause severe socioeconomic damage, accurate high-resolution simulation of precipitation is imperative. However, existing Earth System Models (ESMs) struggle to resolve small-scale dynamics and suffer from biases. Traditional statistical bias correction and downscaling methods fall short in improving spatial structure, while recent deep learning methods lack controllability and suffer from unstable training. Here, we propose a machine learning framework for simultaneous bias correction and downscaling. We first map observational and ESM data to a shared embedding space, where both are unbiased towards each other, and then train a conditional diffusion model to reverse the mapping. Only observational data is used for the training, so that the diffusion model can be employed to correct and downscale any ESM field without need for retraining. Our approach ensures statistical fidelity and preserves spatial patterns larger than a chosen spatial correction scale. We demonstrate that our approach outperforms existing statistical and deep learning methods especially regarding extreme events.
Short Summary: Accurately simulating rainfall is essential to understand the impacts of climate change, especially extreme events such as floods and droughts. Climate models simulate the atmosphere at a coarse resolution and often misrepresent precipitation, leading to biased and overly smooth fields. We improve climate model precipitation using a generative machine learning model that is data-efficient, preserves key climate signals such as trends and variability, and significantly improves the representation of extreme events.
@@ 1 Introduction
With global warming, we anticipate more intense rainfall events and associated natural hazards, e.g., in terms of floods and landslides, in many regions of the world [23]. Understanding and accurately simulating precipitation is particularly important for adaptation planning and, hence, for mitigating damages and reducing risks associated with climate change. Earth System Models (ESMs) play a crucial role in simulating precipitation patterns for both historical and future scenarios. However, these simulations are computationally extremely demanding, primarily because they require solving complex partial differential equations. To manage the computational load, ESMs resort to approximate solutions on discretized grids with coarse spatial resolution (typically around 100 km). The consequence is that these models do not resolve small-scale dynamics, such as many of the processes relevant to precipitation generation. This leads to considerable biases in the ESM fields compared to observations. Moreover, the coarse spatial resolution prevents accurate projections of localized precipitation extremes. Therefore, precipitation fields simulated by ESMs cannot be used directly for impact assessments [44] and especially tasks such as water resource and flood management, which require precise spatial data at high resolution [13].
@@
Statistical bias correction methods can be used as a post-processing to adjust statistical biases. Quantile mapping (QM) is the most common method for improving the statistics of ESM precipitation fields [38, 11, 3, 28]. QM reduces the bias using a mapping that, locally at each grid cell, aligns the estimated cumulative distribution of the model output with the observed precipitation patterns over a reference time period. Although QM is effective in correcting distributions of single grid cells, it falls short in improving the spatial structure and patterns of precipitation simulations [17]. A visual inspection shows that ESM precipitation remains too smooth compared to the observational data after applying quantile mapping.
@@
To address these problems, deep learning methods have recently been introduced [29, 26, 18, 30, 8, 17]. In these approaches, the statistical relationships between model simulations and observational data are learned implicitly. A general constraint when using machine learning methods for bias correction is that individual samples of observational and Earth System Model data are always unpaired. In this context, a sample is a specific weather situation at a specific point in time. The reason for this lack of pairs is that simulations, even with very similar initial conditions, diverge already after a short period of time due to the chaotic nature of the underlying atmospheric dynamics. Currently, one can, therefore, not utilize the wide range of supervised machine learning (ML) techniques that have shown great success in various disciplines in recent years and the available options are consequently restricted to self- and unsupervised machine learning methods. Recent studies [18, 30, 8, 17] applied generative adversarial networks (GANs [10]) and specifically cycleGANs [45] to improve upon existing bias correction techniques. A major limitation of GAN-based approaches is that the stability and convergence of the training process are difficult to control and that it is challenging to find metrics that indicate training convergence. In addition, GANs often suffer from mode collapse, where only a part of the target probability distribution is approximated by the GAN.
@@
As noted above, the low spatial resolution of ESM fields prevents local risk and impact assessment, necessitating the additional use of downscaling methods. In line with the climate literature, we refer to increasing the spatial resolution as downscaling throughout our manuscript, although we are aware that, especially in the machine learning literature, the term upsampling is more prevalent. We use the term downscaling only when we want to increase the information content in an image as well as the number of pixels. When we refer to upsampling (downsampling), we only mean an increase (decrease) in the number of pixels. Statistical downscaling aims to learn a transformation from the low-resolution ESM fields to high-resolution observations. Recent developments lean towards using machine learning methods for this task [31, 22, 32]. The potential for machine learning-based downscaling methods was already shown in [41, 40, 5, 6, 33].
[…]
@@
Diffusion models (DMs) have recently emerged as the state-of-the-art ML approach for conditional image generation [36, 34, 37] and image-to-image translation [35], mostly outperforming GANs across different tasks. Diffusion models (Fig. 1 and fig. S1) avoid the common issues present with GANs in exchange for slower inference speed. A diffusion model consists of a forward and a backward process. During the forward process, noise is added to an image in subsequent steps to gradually remove its content. The amount of noise added follows a predefined equation. During the backward process, a neural network is trained to reverse each of these individual noising steps to recover the original image. The trained diffusion model can generate an image of the training data distribution, given a noise image as input. Recent work [42] introduced a framework for downscaling and bias correction, combining a diffusion model that is responsible for downscaling and a model based on optimal transport responsible for bias correction. Optimal transport [4] learns a map between two data distributions in an unsupervised setting. However, this framework is computationally expensive and has so far only been demonstrated on synthetic datasets, without evaluation on real-world observational or ESM fields. In contrast, our approach is computationally efficient by combining computationally efficient QM for large-scale bias correction with a conditional diffusion model that performs both small-scale bias correction and downscaling by generating matching small-scale patterns. We demonstrate its effectiveness for precipitation data, highlighting its ability to correct biases, downscale accurately, and capture extremes, uncertainties, and trends. A major advantage is that our conditional training allows us to use a relatively small dataset for training and still capture the distribution accurately. In contrast, unconditional models often need considerably more data to capture the full data variability, as we also show in our comparison to [19] (see fig. S2).
Existing work leveraging state-of-the-art ML methods for bias correction and downscaling does not systematically investigate out-of-distribution scenarios like future emission scenarios and especially the representation of extreme events of the generative models in detail. Understanding the generalization performance of the models under these conditions is, however, crucial for impact modelers who rely on these outputs for risk assessments under future climate conditions. We will therefore present a detailed analysis of the generalization capability of our approach, both in terms of its performance in preserving climate change trends, as well as in capturing extreme events and their trends.
A major challenge in bias correction and downscaling of ESMs is that the whole class of state-of-the-art supervised machine learning methods is not applicable in this setting. This is due to two fundamental issues. First, due to the chaotic nature of atmosphere and ocean dynamics, ESM simulations and observational data are inherently unpaired. This means that the weather on a specific day in an ESM simulation does not correspond to the observed weather on the same day, which prevents directly training a supervised ML method on the task. Second, training a ML model on observational data and applying it to ESM data is unreliable due to the substantial distribution shift between both datasets caused by systematic biases in the ESM. This violation of the assumption of independently and identically distributed (i.i.d.) data leads to poor generalization. Our proposed framework directly addresses both challenges. We reformulate the problem in a novel way, which allows us to train arbitrary ML models in a conditional setup without the need for explicit ESM-observation pairs, while at the same time resolving the distribution shift.
We present a novel framework based on state-of-the-art conditional diffusion models that allows us to perform both bias correction and downscaling with one single neural network, which only takes precipitation as input and output. We use a conditional diffusion model (Fig. 1 and fig.  S1) to correct low-resolution (LR) ESM fields toward high-resolution (HR) observational data (OBS). The supervised formulation of the task allows us to train a conditional diffusion model that is more data efficient (requiring less training data) than its unconditional counterpart because it is trained to only learn the small-scale precipitation patterns, given the large-scale patterns. The model then learns to copy the correct large-scale information from the condition channel. An unconditional model that learns to approximate the full distribution of precipitation at all scales is unnecessarily complex for the task. In general, our task of bias correction and downscaling can be seen as taking a field from a distribution $p(\text{ESM})$ and transforming it into a field from a conditional distribution $p(\text{OBS}|\text{ESM})$.
A key idea of our framework is to reformulate the problem in a way that yields a clear training objective. A key part of it is a statistical mapping to an embedding space, which ensures that training and inference data are identically distributed. We achieve this by introducing transformations $f$ and $g$ that map observational (OBS) and ESM data to a shared embedding space (see Methods and Fig. 1A). This space is explicitly designed to solve the two fundamental issues mentioned above: it creates a valid supervised objective by providing paired samples of observational data and their perturbed embeddings (OBS, f(OBS)), and it ensures the training and inference distributions match by making the distributions of the embeddings $f(OBS)$ and $g(ESM)$ similar. On this shared embedding space, we can train a conditional diffusion model to approximate the inverse of $f$ (Fig. 1B and Fig. 1C). The neural network is trained to predict the clean OBS data given the embedded OBS data, thereby only relying on pairs between OBS and $f(OBS)$. For inference, the ESM data is mapped into the same embedding space using the transformation $g$. The statistical similarity of the resulting embeddings $f(OBS)$ and $g(ESM)$ enables the diffusion model, which was trained exclusively on observational data, to generalize effectively to downscale and bias-correct the ESM fields. The diffusion model will map the embedded ESM data towards the distribution of observational data, resulting in bias-corrected and downscaled ESM fields.
This framework offers great flexibility as it can be applied to any ESM, with minimal adjustments in the embedding pipeline. The embedding transformation for the ESM has two key components. First, we use quantile mapping (QM) as a fast and effective method to correct large-scale biases in the ESM. Second, we introduce noise to remove small-scale information in the precipitation fields. We define large scales as those spatial scales that are effectively corrected using QM alone, while smaller spatial scales, which require additional correction, are referred to as small scales (Fig. 2). This noise selectively targets small-scale patterns, leaving intact large-scale patterns. In our approach, quantile mapping addresses large-scale biases, while the small-scale biases and downscaling are handled by our diffusion model. The task of our model is then to perform downscaling and bias correction by regenerating these small-scale features, in a way that ensures consistency with the preserved large-scale patterns. When applying our framework on a different region or ESM, it is computationally inexpensive to recompute the quantile mapping (QM) for the embedding transformation.
@@
Figure 1: Schematic overview of our approach. (A) Bias correction and downscaling can be formulated as a mapping $\omega$ from the ESM data space to the data space of observations (OBS) used for training. We first map both datasets to a shared embedding space and then learn the inverse of the mapping $f$ with a DM. We achieve a correction of the ESM data by applying $DM\circ g$. (B) Our framework allows to train a single model for bias correction and downscaling in a supervised way despite the unpaired nature of OBS and ESM fields. We construct functions $f,g$ that map $OBS\in\mathbf{V^{obs}}$ and $ESM\in\mathbf{V^{esm}}$ fields to a shared embedding space $\mathbf{V^{emb}}$. Note that this embedding space does not enforce pairing between individual fields, but a similar distribution between the embedded fields. By inverting $f$, we can rewrite $\omega$ as $\omega=f^{-1}\circ g$. We learn the inverse $f^{-1}$ with a conditional diffusion model. This model is trained (blue arrow) on pairs of observational data to approximate the map from $f(OBS)$ to OBS. Because $f(OBS)$ and $g(ESM)$ share the embedding space (and are identically distributed by construction), we can evaluate (green arrow) the DMDM on the embedded ESM data $g(ESM)$ and thereby approximate the bias correction and downscaling function $\omega=f^{-1}\circ g\approx DM\circ g$, without the need of paired data between OBS and ESM. The indices $i,j$ highlight that the two exemplary fields $ESM_{j}$ and $ESM_{i}$ are not paired.
@@
(C) Left: Training process of the conditional DM $DM\approx f^{-1}$. Note that the individual samples of the input OBS and their embeddings $f(OBS)$, as well as the embeddings $f(OBS)$ and the output of $DM\approx f^{-1}$ are paired, respectively. Right: Inference process of $DM\approx f^{-1}$. In this case, the individual samples of the input ESM, their embeddings $g(ESM)$, and the output of $DM\approx f^{-1}$ are paired, respectively. It is not necessary for the training embedding samples to be paired with the inference embedding samples. See fig. S1 for details.
@@ 2 Results
Figure 2: Power spectral densities (PSDs) for different choices noising scale of the diffusion model. The noising scale $s$ (dashed line) is a hyperparameter that can be chosen depending on the ESM and observational datasets, as well as on the specific task. For the maximal choice of $s$ (A) all information in the observations (ERA5) and model simulations (GFDL) is noised and thereby destroyed. Conditioning on pure noise makes the task equivalent to unconditional image generation. The diffusion model will learn to generate observational fields with no relation to the ESM fields. When $s$ is chosen to be minimal, there will be no noising and the conditional generation will directly replicate the condition, i.e. the ESM field. In (B) we chose $s$ as the point where the PSDs of the observational and simulated datasets intersect. We then apply sufficiently many forward diffusion noising steps to both datasets, destroying small-scale structure until they agree in the PSD. We call scales smaller than $s$ small scales and scales larger than $s$ large scales. In (C) and (D), the effects of choosing a smaller noise scale $s$ are shown. Prior knowledge about the ESM or its accuracy can also guide the choice of $s$.
@@ 2.1 Embedding evaluation
Transformations $f,g$ are chosen so that they map observational (OBS) and model (ESM) data to a common embedding space $\mathbf{V^{emb}}$, where all samples are identically distributed. For constructing $f$ and $g$ we need $f(ERA5)$ and $g(GFDL)$ to be unbiased with respect to each other. The transformations need to be chosen such that the embedded data share the same distribution and the same power spectral density (PSD). We assess if they are statistically unbiased towards each other by analyzing their histograms and latitude / longitude profiles, as well as their spatial PSDs (after applying pre-processing transformations). Figure S3 shows that $f(ERA5)$ and $g(GFDL)$ have the same spatial distribution (fig. S3A) with minor differences in temporal statistics shown by the histogram (fig. S3B) and latitudinal/ longitudinal profiles (fig. S3C and fig. S3D).
@@
The individual operations that make up the transformations $f$ and $g$ do not change the large-scale patterns of their respective inputs, as desired for a valid bias correction. The goal of downscaling and bias correction $\omega$ (Fig. 1) is to rely on the unbiased large-scale patterns of the ESM and correct statistics, as well as small-scale patterns. The transformation $g$ preserves the unbiased information from the ESM by construction. Therefore, we want the diffusion model, approximating $f^{-1}$, to also preserve unbiased information.
@@
The extreme case of erasing all detail with large amounts of noise (Fig. 2A) leads to learning the unconditional distribution $p(ERA5)$, which is then not a correction of $GFDL$ but a generative emulation of the ERA5 reanalysis data. We tested this by adding the same amount of noise to the output of our diffusion model that was added to create $g(GFDL)$. This ensures that both the downscaled and bias-corrected fields, as well as the original GFDL fields, lack the small-scale details up to the same point.
@@
To verify that large-scale patterns are preserved by the diffusion model, we compute image similarity metrics between the low pass filtered version of the input of the diffusion model (embedded ERA5 data $f(ERA5)$) and the low pass filtered output of the diffusion model $DM(f(ERA5))$. The output of the low pass filter leaves the large-scale features unchanged. The comparison yields an average structural similarity index (SSIM [43]) value of 0.85 and a Pearson correlation coefficient of 0.95 for the validation dataset. This verifies that large-scale patterns are well preserved by the diffusion model.
@@
Our diffusion model is able to reconstruct high-resolution fields following the ERA5 distribution from embedded ERA5 fields $f(ERA5)$, with only minor discrepancies in small-scale patterns (fig.  S4A). A comparison between the mean absolute spatial-temporal difference between the first downsampled and then bilinearly upsampled ERA5 and the ground truth ERA5 fields at 0.25° yields a mean bias of 0.27 mm d-1. The downscaling of our diffusion model reduces this bias to 0.21 mm d-1 (at 0.25°). Our diffusion model thus approximates $f^{-1}$ well, and we successfully created a shared embedding space in which $f(ERA5)$ and $g(GFDL)$ are identically distributed.